### v2

# When higher-order interactions enhance synchronization: the case of the Kuramoto model on random hypergraphs

Abstract
Synchronization is a fundamental phenomenon in complex systems, observed across a wide range of natural and engineered contexts. The Kuramoto model provides a foundational framework for understanding synchronization among coupled oscillators, traditionally assuming pairwise interactions. However, many real-world systems exhibit group and many-body interactions, which can be effectively modeled through hypergraphs. Previous studies suggest that higher-order interactions shrink the attraction basin of the synchronous state, making it harder to reach and potentially impairing synchronization, despite enriching the dynamics. In this work, we show that this is not always the case. Through a numerical study of higher-order Kuramoto models on random hypergraphs, we find that while strong higher-order interactions do generally work against synchronization, weak higher-order interactions can actually enhance it when combined with pairwise ones. This result is further corroborated by a cost-benefit analysis: under a constrained budget of both pairwise and higher-order interactions, a mixed allocation involving both consistently achieves higher synchronization than relying on either interaction type alone. These findings provide new insights into the role of higher-order interactions in shaping collective dynamics and point to design principles for optimizing synchronization in complex systems.
Introduction
Synchronization dynamics is one of the most studied phenomena in the field of complex systems and a paradigmatic example of self-organizing behavior Pikovsky et al. (2001); Arenas et al. (2008); Boccaletti et al. (2018). The synchronization capabilities of coupled oscillators were first observed by Christiaan Huygens, who, in the 17th century, noticed that pendula hanged on the same wall tend to synchronize in anti-phase. Later research highlighted the occurrence and relevance of synchronization in many natural and artificial systems, ranging from fireflies’ blinking and frogs’ croaking to rhythmic contraction in cardiac cells and neuronal activity, and from bridge oscillations to power angles in power grids Strogatz (2003).
The key factor enabling this collective behavior are the interactions among the oscillators, which can be, e.g., mechanical (bridges, pendulum clocks), electrical (heart, power grids), visual (fireflies), or acoustic (frogs) Pikovsky et al. (2001); Strogatz (2003); Coraggio et al. (2025). Indeed, first Winfree understood that such interactions could be thought as a perturbation Winfree (1967), then Kuramoto Kuramoto (1975) obtained a simple model in which synchronization emerges as a phase transition triggered by the variation in the coupling strength: the celebrated Kuramoto model Strogatz (2000); Acebrón et al. (2005). In the classic Kuramoto model, two assumptions are taken which do not necessarily hold in applications. First, interactions are assumed to be pairwise, i.e., each one involving exactly two oscillators. Second, an all-to-all interaction topology is assumed, where each oscillator interacts with all others with the same strength.
However, research has shown that, when different complex topologies are considered, the dynamics of the Kuramoto model is much richer, displaying complex patterns and transitions Rodrigues et al. (2016), generally making synchronization more difficult to achieve. This shift is motivated by the applications, where all-to-all topologies are actually rare, and the interactions are modeled through networks Newman (2018); Latora et al. (2017). Despite its versatility, the network approach has the limitation of considering only pairwise interactions. Indeed, many-body (rather than pairwise) couplings have been found to be better suited to describe certain kinds of interactions, such as those occurring in social sciences, ecology, and neuroscience Battiston et al. (2020); Bianconi (2021); Battiston et al. (2021); Bick et al. (2023); Boccaletti et al. (2023); Muolo et al. (2024a); Millán et al. (2025). Mathematically, such many-body interactions can be modeled via hypergraphs and simplicial complexes, which are extensions of networks (i.e., graphs).
Over the past few years, it has been shown that higher-order interactions enrich system dynamics, with notable applications in the synchronization of chaotic oscillators Krawiecki (2014); Gambuzza et al. (2021); Gallo et al. (2022); Della Rossa et al. (2023), chimera states Kundu and Ghosh (2022); Muolo et al. (2024b), random walks Schaub et al. (2020); Carletti et al. (2020a), pattern formation Carletti et al. (2020b); Muolo et al. (2023), opinion dynamics Iacopini et al. (2019); Neuhäuser et al. (2020); DeVille (2021), and pinning control De Lellis et al. (2022); Xia and Xiang (2024); Muolo et al. (2025), to name a few. In particular, extensive literature shows that higher-order interactions enrich the dynamics of the Kuramoto model, generally making synchronization more difficult to achieve Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020); Lucas et al. (2020); Adhikari et al. (2023); Skardal et al. (2023); León et al. (2024); Costa et al. (2024); Huh and Kim (2024); Wang et al. (2024); Zhang et al. (2024). In particular, Zhang et al. Zhang et al. (2024) showed that higher-order interactions make the attraction basin of the synchronous state smaller but more robust (deeper). These findings were further supported in Fariello and de Aguiar (2024), showing that the critical coupling for synchronization increases, facilitating desynchronization, and in von der Gracht et al. (2024); Wang et al. (2025a), where it was demonstrated that, once achieved, synchronization and twisted states becomes harder to disrupt due to higher-order interactions. Overall, while it is well known that stronger pairwise interactions enhance the degree of synchronization of phase oscillators in the purely pairwise case Kuramoto (1984); Acebrón et al. (2005); Rodrigues et al. (2016), the role of higher-order interactions remains far from fully understood.
These recent findings motivate our first research goal, which is to quantify the effects of higher-order interactions on the synchronous state for more general hypergraph topologies. As in previous works Tanaka and Aoyagi (2011); Bick et al. (2016); Skardal and Arenas (2019, 2020); Millán et al. (2020); Lucas et al. (2020); Adhikari et al. (2023); Skardal et al. (2023); León et al. (2024); Costa et al. (2024); Huh and Kim (2024); Wang et al. (2024); Zhang et al. (2024); Fariello and de Aguiar (2024); von der Gracht et al. (2024); Wang et al. (2025a), our framework considers the addition of higher-order interactions to the pairwise setting. To this end, we studied Kuramoto oscillators coupled through $2$- and $3$-body interactions and conducted a numerical study on random hypergraphs of $10$ to $100$ nodes, a regime relevant for real-world applications (e.g., Menara et al. (2022); Simpson-Porco et al. (2013); Alderisio et al. (2017)). Our analysis confirms that, as expected, higher-order interactions generally enhance synchronization when the initial conditions are close to the synchronous state; however, when starting from incoherent states (i.e., far from synchronization), things become more interesting. While our simulations support the finding that higher-order interactions shrink the attraction basin of the synchronous state, they also reveal that weak higher-order interactions can enhance the degree of synchronization, when added to pairwise ones. As Zhang et al. Zhang et al. (2024) cleverly noted, the attraction basin becomes “deeper but smaller” with higher-order interactions. Our results seem to corroborate this observation and may be interpreted through the observation that the basin deepens before it begins to shrink.
Our second research question is: given a limited amount of resources for connectivity of both pairwise and higher-order interactions, which is the optimal combination to enhance synchronization? With this investigation, we aim to determine whether higher-order interactions can offer advantages over purely pairwise ones, and whether a mix of both can outperform structures relying exclusively on one type. This question is relevant in both engineered and natural systems: in the former, it can guide resource allocation for building synchronizable systems; in the latter, it may help explain the interaction patterns that emerge in nature as evolved or self-organized solutions to synchronization demands.
While the problem of optimally allocating links to promote specific collective behaviors, and in particular synchronization, has been extensively studied in complex networks, only a few studies have explored the role of higher-order interactions Lamata-Otín et al. (2025). Much work focused on identifying metrics that promote or inhibit synchronization, showing, e.g., that factors contributing to good synchronizability are low eigenratio Pecora and Carroll (1998), high algebraic connectivity Coraggio et al. (2018), or high minimum density Coraggio et al. (2021, 2020). General network structural features also play a critical role: for example, small-world networks synchronize more effectively than random graphs Barahona and Pecora (2002). Notably, Donetti et al. Donetti et al. (2005, 2006) observed that optimally synchronizable networks tend to exhibit “entangled” homogeneous structures, while Skardal et al. Skardal et al. (2014) found that matching between frequency heterogeneity and network heterogeneity enhances synchronization. Several approaches have also been proposed, from a network synthesis perspective, to optimally design synchronizable networks, both for oscillators ensembles Fazlyab et al. (2017); Lei et al. (2023) and for general dynamical systems Nishikawa and Motter (2006); Estrada et al. (2010); see also Coraggio et al. (2026) and references therein. Importantly, it has been shown that optimal network topologies can vary significantly when individual nodes dynamics are considered, highlighting a complex interplay between structure and dynamics, and, consequently, the coupling laws governing the interactions Gorochowski et al. (2010); Coraggio and di Bernardo (2024).
Many of the mentioned studies focus on chaotic oscillators, whose synchronization is studied through the Master Stability Function, a linear stability analysis technique first proposed by Fujisaka and Yamada Fujisaka and Yamada (1983) and later extended to general networks by Pecora and Carroll Pecora and Carroll (1998), who also coined its current name. This approach, however, is not suited to coupled phase oscillators such as the Kuramoto model. On the other hand, the method exploited in Coraggio and di Bernardo (2024), can be also applied to phase oscillators and partly inspired the numerical approach we employed to answer this question. Our analysis on random hypergraphs shows that, when the total budget for interactions (pairwise and higher-order) is limited, synchronization is enhanced by a combination of both types of interactions, regardless of the relative cost of higher-order interactions within that budget.
The two main findings of this work—that (i) weak higher-order interactions enhance synchronization (i.e., increase the order parameter on average), when added to pairwise ones and that (ii) with a finite budget for connections, a combination of pairwise and higher-order interactions optimizes synchronization—were obtained via a numerical analysis of the higher-order Kuramoto model on random hypergraphs, and are summarized in Figure 1. These results are compatible with previous evidence for the all-to-all case, where it was shown that synchronization increases when higher-order interactions are added to pairwise ones Skardal and Arenas (2020); Kuehn and Bick (2021); Chen et al. (2025), and with those of two recent independent preprints, which have partially corroborated and complemented our settings by focusing on hyperrings Wang et al. (2025b); Skardal et al. (2025).
Figure 1: a) Meaning of the order parameter $R$: high values correspond to synchronized states, whereas low values indicate incoherent dynamics. b) Pictorial representation of the main findings of this work. (1) While higher-order interactions generally hamper synchronization by making the attraction basin of the synchronous state “deeper but smaller”, adding weak higher-order interactions to pairwise-coupled networks enhances synchronization. (2) Under a finite budget for interactions, regardless of the relative cost of higher-order interactions, the optimal configuration for synchronization always involves a combination of pairwise and higher-order interactions. Red (resp. pink) hyperedges denote higher-order interactions with larger (resp. smaller) coupling strength. Yellow (resp. blue) nodes indicate highly (resp. weakly) synchronized states, in accordance with panel a).
In the next section, we present the results of the two numerical studies on the higher-order Kuramoto model with $3$-body interactions. The functional form of such interactions is not unique; in the Main Text, we show results for one form, while results for a different one, show a qualitatively similar behavior and are included in the Supplementary Materials. In the Methods section, we provide rigorous definitions of the framework, detail the derivation of the higher-order Kuramoto model and the choice of interaction terms, and describe the numerical methods used.
Results
To contextualize our results, we first briefly introduce the Kuramoto model with higher-order interactions that we considered in this study. We analyze a system of $N_{0}$ phase oscillators. Here, we present results for $N_{0}=10$; simulations for $N_{0}=100$—reported in the Supplementary Materials—exhibit the same qualitative behavior. The interactions are encoded by the adjacency matrix $A^{(1)}\in\{0,1\}^{N_{0}\times N_{0}}$ (i.e., the first order adjacency tensor) and the second order adjacency tensor $A^{(2)}\in\{0,1\}^{N_{0}\times N_{0}\times N_{0}}$, respectively. We assume undirected, unweighted, and unsigned interactions; i.e., connections are either present or absent.
In particular, $A^{(1)}_{jk}=1$ if oscillators $j$ and $k$ are connected by a $2$-body interaction (a link), and 0 otherwise; $A^{(2)}_{jkl}=1$ if oscillators $j$, $k$, and $l$ are connected by a $3$-body interaction (a triangle), and 0 otherwise [^1] In the literature, a $2$-body interaction is equivalently referred to as a link, edge, pairwise interaction, $1$-hyperedge, or $1$-simplex. A $3$-body interaction is also called a triangle, $2$-hyperedge, or $2$-simplex. (see Methods for further details). The oscillators are assigned natural frequencies $\omega_{j}$, for $j\in\{1,...,N_{0}\}$, drawn randomly from a uniform distribution of a given interval. Simulations with normally distributed frequencies are reported in the Supplementary Materials: we observed no qualitative difference between the two cases. Hence, the higher-order Kuramoto model we consider is governed by the following equations, for $j\in\{1,\dots,N_{0}\}$:
$\dot{\theta_{j}}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})+\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}),$ (1)
where $\langle d^{(1)}\rangle$ and $\langle d^{(2)}\rangle$ are the average $1$- and $2$-degree of $2$- and $3$-body interactions, respectively, and $K_{1}$ and $K_{2}$ are the coupling strengths of the $2$- and $3$-body interactions, respectively. Following Namura et al. (2025), we refer to the higher-order interaction term $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ used in (1) as the (1,1,2)(1,1,-2) interaction. Results for a different $3$-body interaction, namely $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$, which we call the (2,1,1)(2,-1,-1) interaction, are reported in the Supplementary Materials, and show a qualitatively similar behavior to the simulations presented here. The level of synchronization is measured by the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\theta_{j}}\right\rvert$.
Effects of coupling strengths on synchronization
To further validate the claim that the attraction basin of the synchronous state becomes “deeper but smaller” Zhang et al. (2024) as the strength of higher-order interactions increases (in the presence of pairwise interactions), we conducted a systematic study of the higher-order Kuramoto model.
Specifically, we first fix the total number of nodes ($N_{0}$), links ($N_{1}$) and triangles ($N_{2}$) and then generate $N_{\mathcal{H}}$ connected hypergraphs randomly allocating all links and triangles. We consider two connectivity conditions: $\mathcal{H}$-connectivity, if the projected network is connected, and $1$-connectivity, if the structure is already connected solely with the links. These two types of connectivity are illustrated in Figure 2, and their rigorous definitions are provided in the Methods section. Then, we ran the simulations varying the coupling strengths of pairwise and higher-order interactions, $K_{1}$ and $K_{2}$. For each pair, we performed $N_{\mathcal{H}}=300$ integrations over random hypergraphs of $N_{0}=10$ nodes, initial conditions and frequency distributions, averaging the resulting order parameter $R$ (see Methods for details). Simulations for $N_{\mathcal{H}}=100$ integrations over random hypergraphs of $N_{0}=100$ nodes give analogous results, which are displayed in the Supplementary Materials.
Figure 2: a) Examples of $\mathcal{H}$-connected hypergraphs, i.e., whose projected networks (bottom row) are connected. b) Examples of $1$-connected hypergraphs, i.e., whose underlying networks are already connected.
Figure 3: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order 33-body interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs, using the (1,1,2)(1,1,-2) interaction. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$: that value is never zero. Top panels show the case of $\mathcal{H}$-connected hypergraphs, while bottom panels show the case of $1$-connected hypergraphs: we observe no significant differences between the two.
Adding weak higher-order interactions enhances synchronization
The top panels show the case of $\mathcal{H}$-connected hypergraphs, while the bottom ones portray the case of $1$-connected hypergraphs. We see no significant differences, indicating that the observed effects arise from higher-order interactions, rather than connectivity differences. The left panels show the case in which the system starts close to the synchronized state. As expected from previous studies Zhang et al. (2024); von der Gracht et al. (2024), higher-order interactions enhance synchronization, as evidenced by the yellow regions (i.e., higher values of the order parameter $R$) widening as $K_{2}$ increases. The right panels, instead, show the case of incoherent initial conditions. Again, as expected Zhang et al. (2024); Fariello and de Aguiar (2024); Muolo et al. (2025), we see that the stronger higher-order interactions ($K_{2}$) worsen synchronization (lower $R$). However, for small $K_{2}$ (weak higher-order interactions) $R$ initially increases before eventually declining. To better visualize this effect, we computed, for each $K_{1}$, the maximum value of $R$ in the case where the system starts from incoherence (right panels, black line).
Since weak higher-order interactions lead to higher values of $R$ also when starting from coherent initial conditions (left panels), we conclude that weak higher-order interactions enhance synchronization on average, when pairwise ones are already present.
We also tested $K_{1}$ and $K_{2}$ values up to $1.5$ (results omitted here for brevity), confirming the same trend.
In the Supplementary Materials, we report a figure analogous to Figure 3, obtained assessing the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$, along with results for the (2,1,1)(2,-1,-1) interaction for $3$-body interactions, normally distributed frequencies, and larger hypergraphs. All findings are consistent with those reported in the Main Text.
Emergence of distinct dynamical regimes
As shown in previous works Tanaka and Aoyagi (2011); Skardal and Arenas (2020); León et al. (2024); Battiston et al. (2025), higher-order interactions can induce bistability [^2] Coexistence of locally stable synchronized states and locally stable incoherent states. in the Kuramoto model. Consistently, Figure 3 shows a behavior compatible with bistability for low $K_{1}$ and high $K_{2}$: trajectories starting near the synchronous state remain synchronized, whereas incoherent initial conditions do not typically lead to synchronization.
To systematically explore the dependence of the system’s behavior on the coupling strengths $K_{1}$ and $K_{2}$, Figure 4 reports the distributions of the order parameter $R$ for representative parameter values across different realizations of hypergraphs, initial conditions, and frequency distributions (bottom panels), focusing on initially incoherent $1$-connected hypergraphs (Fig. 3, bottom-right panel). The top-left and top-right panels of Figure 4 show bifurcation diagrams of the average $R$ (blue line), together with its standard deviation (shaded area), obtained by varying one coupling strength while keeping the other fixed. An analogous analysis for Gaussian frequency distributions is provided in the Supplementary Materials.
For fixed $K_{1}=0.15$, large $K_{2}$ (pink upward triangle marker) yields a distribution of $R$ spanning both low and high values, which is compatible with bistability Suman and Jalan (2024). In contrast, for smaller $K_{2}$ (gray circle and pink downward triangle), despite a similar mean value of $R$, the distribution is more concentrated around the mean, indicating a more consistent behavior across realizations. For fixed $K_{2}=0.25$, increasing $K_{1}$ leads to distributions of $R$ that become increasingly concentrated near $1$, indicating a progressive stabilization of strongly coherent states across realizations.
Figure 4: System behavior with different pairwise ($K_{1}$) and higher-order ($K_{2}$) coupling strengths, for initially incoherent $1$-connected hypergraphs from Fig. 3. The top-center panel replicates the bottom right panel of Fig. 3 for comparison (color is the mean order parameter $R$). The top-left (resp. top-right) panel shows the mean value of $R$ (blue line) and its standard deviation (shaded area) as a function of the higher-order coupling strength $K_{2}$ (resp. $K_{1}$) while keeping the pairwise interaction strength $K_{1}$ (resp. $K_{2}$) fixed. The values of $K_{1}$, $K_{2}$ explored are also indicated in the top-center panel by a horizontal (resp. vertical) black line. Pink (resp. brown) triangles pointing upward/downward (resp. leftward/rightward) mark representative pairs $(K_{1},K_{2})$. The bottom panels display the distributions (integral normalized to 11) of $R$ across the $N_{\mathcal{H}}=300$ realizations of hypergraphs, initial conditions, and frequencies. For fixed $K_{1}=0.15$, a small $K_{2}$ (pink downward triangle) produces $R$ values relatively concentrated around the mean, whereas larger $K_{2}$ (pink upward triangle) leads to a broader distribution, spanning incoherent and strongly coherent regimes. For fixed $K_{2}=0.25$, increasing $K_{1}$ (brown leftward and upward triangles) causes the distribution to concentrate near $1$, indicating a consistently high level of synchronization.
Optimal hyperedge allocation for synchronization
Figure 5: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the (1,1,2)(1,1,-2) interaction. $K_{1}$ and $K_{2}$ are the coupling strengths of $2$- and $3$-body interactions. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color). See Methods for further details. We observe that combining links and triangles generally yields higher synchronization than using either alone.
Next, we explore how limited resources allocated between pairwise and higher-order interactions influence synchronization. To model resource constraints, we assume a fixed budget of $J=40$ arbitrary units, representing energy, material, or financial cost. Each link incurs a cost of $1$ arbitrary unit, while triangles cost $c_{2}$ units, with $c_{2}\in\{1,3,5\}$. Resources are allocated proportionally: a fraction $b_{1}$ of the budget is dedicated to links, while the remaining fraction $1-b_{1}$ is allocated to triangles: specifically, the numbers of links and triangles are $N_{1}=\lfloor Jb_{1}/c_{1}\rfloor$ and $N_{2}=\lfloor J(1-b_{1})/c_{2}\rfloor$, respectively; both links and triangles are then chosen randomly, ensuring the resulting hypergraph is $\mathcal{H}$-connected. This allocation models practical scenarios where different interactions types carry heterogeneous resource costs. Beyond structural costs, synchronization also depends on coupling strengths $K_{1}$, $K_{2}$—associated with links and triangles, respectively—which are typically determined by the system’s physics. Our goal is to identify the optimal allocation strategy across varying relative costs and a range of coupling strength configurations. Additionally, we investigated how the coherence of initial conditions—from partially aligned to fully disordered—impacts the effectiveness of different allocation strategies. To systematically explore these factors, we evaluated synchronization across a grid of $51\times 51=2{,}601$ experimental conditions, varying coupling strength distributions ($K_{1}$, $K_{2}$) and allocation ratios ($b_{1}$). Synchronization performance was quantified by the time-averaged Kuramoto order parameter $R$, averaged over $N_{\mathcal{H}}=500$ random realizations per experimental condition to capture variability in structure, frequencies, and initial states. This exploration was repeated $6$ times in total, varying both the relative cost of triangles ($c_{2}\in\{1,3,5\}$) and the coherence of initial conditions ($\theta_{0}^{\mathrm{max}}=\pi/2$, for more coherent starts and $\theta_{0}^{\mathrm{max}}=2\pi$ for fully incoherent starts). Figure 5 summarizes the results of this extensive simulation campaign.
Within each panel, the horizontal axis represents the fraction of resources allocated to links, $b_{1}$, and the vertical axis represents their associated coupling strength, $K_{1}$, with the coupling strength for triangles set as $K_{2}=0.3-K_{1}$; the sum of the two strengths is kept constant to ensure meaningful comparison across conditions. The color of each pixel denotes the order parameter $R$, averaged over both the second half of simulations and the $500$ hypergraph realizations, with brighter colors indicating stronger phase coherence. The black line denote the optimal link allocation fraction $b_{1}$ for each combination of coupling strengths $K_{1}$ and $K_{2}$.
Combining links and triangles is optimal
A key finding from this analysis is that the optimal link allocation fraction $b_{1}$ is rarely exactly 0 or $1$, i.e., dedicating all resources exclusively to either links or triangles is almost never optimal. Trivially, full allocation to links is optimal only when $K_{2}\approx 0$; similarly, allocating the entire budget to triangles is optimal only when $K_{1}\approx 0$. Across all panels, the optimal configuration consistently favors a combination of both interaction types. Notably, the optimal value of $b_{1}$ strongly depends on the relative coupling strengths $K_{1}$, $K_{2}$, with more resources typically allocated to the interaction type associated with the stronger coupling.
Triangles are less effective but still useful with incoherent initial conditions
Surprisingly, the beneficial effect of triangles persists even when the initial phases are highly incoherent (right panels). In this scenario, triangles are less effective at enhancing synchronization (see the lower-left corners of the right panels). Nevertheless, a combination of interaction types remains optimal, as highlighted by the black lines indicating the best $b_{1}$ for each pair $(K_{1}$, $K_{2})$. In all cases, including some triangles improves synchronization, with the exact optimal balance depending on the relative values of the coupling strengths $K_{1}$ and $K_{2}$.
Triangles remain beneficial even when more costly
As expected, lower costs for triangles lead to generally higher levels of synchronization (compare the left panels, from top to bottom). This is because, with a fixed resource budget, a lower cost allows deploying more triangles, resulting in a more connected hypergraph. Interestingly, increasing the relative cost $c_{2}$ of triangles shifts the optimal resource allocation: despite the higher cost, a larger portion of the budget is devoted to triangles, underscoring their strong impact on synchronization. This is reflected by the black optimal-allocation curves bending slightly left as $c_{2}$ increases.
These findings indicate that higher-order interactions complement pairwise couplings and can enhance synchronization. To further validate this, we repeated the analysis under different conditions: using the (211)(2-1-1) interaction as a different coupling function for the $3$-body interactions Skardal and Arenas (2020), inspecting the Kuramoto-Daido order parameter $R_{2}$ instead of the Kuramoto order parameter $R$, drawing the oscillator frequencies randomly from a Gaussian distribution rather than a uniform one, and considering larger hyperegraphs of $N_{0}=100$ nodes. In all cases, the results consistently confirm that a combination of links and triangles is more effective for synchronization than relying on a single interaction type. These additional results are reported in the Supplementary Materials.
Discussion
Recent work on nonlinear oscillatory systems has shown that higher-order interactions often lead to richer dynamics and generally hinder synchronization. Except when the system starts within the attraction basin of the synchronous state Zhang et al. (2024); von der Gracht et al. (2024), which higher-order interactions make more robust, the presence of higher-order interactions generally conspires against synchronization, while other dynamical patterns occur, such as cluster synchronization, slow-switching, and multi-stability Tanaka and Aoyagi (2011); Skardal and Arenas (2020); León et al. (2024).
In this work, we performed a numerical study of the higher-order Kuramoto model on random hypergraphs—taken as a paradigmatic example of oscillatory dynamics with higher-order interactions—spanning various parameter configurations. We confirmed that higher-order interactions make the synchronous state more robust. At the same time, while strong higher-order interactions were found to hamper synchronization, weak higher-order interactions can instead enhance it (i.e., increase the order parameter, on average), when added to pairwise ones. This conclusion was further supported by a cost analysis, which showed that a combination of both pairwise and higher-order (3-body) interactions yields the most effective topology for achieving synchronization. Since our results are obtained by averaging over hundreds of random hypergraphs, they hold on average; specific hypergraph structures may behave differently, as reflected in the order parameter distributions in Figure 4.
These findings are consistent with previous analytical results obtained in all-to-all settings, both for the (1,1,2)(1,1,-2) interaction Chen et al. (2025) and the (2,1,1)(2,-1,-1) one Skardal and Arenas (2020), where it was shown that synchronization is slightly enhanced for weak higher-order coupling, prior to the possible onset of bistability at larger coupling strengths, or of explosive phenomena in general Kuehn and Bick (2021). Related results have also been reported independently in recent preprints on hyperring topologies, showing that weak higher-order interactions Wang et al. (2025b), as well as combinations of pairwise and higher-order interactions Skardal et al. (2025), enhance the stability of twisted states while preserving the structure of the attraction basin. Taken together, our results and those reported in the literature complement the picture introduced by Zhang et al. Zhang et al. (2024), according to which higher-order interactions make attraction basins “deeper but smaller”, by showing that weak higher-order interactions first affect basin depth before influencing basin size.
Due to the computational cost of our analyses, we focused on random hypergraphs with up to $100$ nodes, which are, nonetheless, relevant for real-world applications (e.g. Menara et al. (2022); Simpson-Porco et al. (2013); Alderisio et al. (2017)). In light of previous analytical results for all-to-all configurations Skardal and Arenas (2020); Chen et al. (2025), it is plausible that the phenomena identified here persist in larger systems with link and triangle densities comparable to or higher than those considered in this work, whereas sparser configurations may require ad hoc investigations.
An in depth exploration of different higher-order topologies represents a natural direction for future work. Another promising avenue would be to identify which higher-order topologies most effectively enhance synchronization, by extending the numerical network analysis in Coraggio and di Bernardo (2024) to the higher-order setting. Analytical studies on general hypergraph topologies are also needed to develop a comprehensive theoretical framework, to identify possible counterexamples, and to assess whether different frequency distributions (and connectivity type, e.g., $1$-connected vs. $\mathcal{H}$-connected) affect these phenomena in the thermodynamic limit and in very-large-scale systems, as frequency distributions are known to markedly shape the dynamics Pazó (2005). While this remains an open problem, the self-consistency analysis Ichinomiya (2004); Restrepo et al. (2005) developed by Chen et al. for the all-to-all higher-order setting Chen et al. (2025) provides a promising starting point.
In conclusion, the insights gained in this study offer guidance for the design and control of complex systems with higher-order interactions D’Souza et al. (2023). We expect these results to stimulate further theoretical investigations and to contribute to a deeper understanding of the interplay between higher-order topology and nonlinear dynamics at a fundamental level.
Methods
Hypergraphs: basic definitions
We consider higher-order Kuramoto models, with interactions described by undirected, unweighted, unsigned hypergraphs. A hypergraph is defined as $\mathcal{H}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{1,\dots,N_{0}\}$ is the set of $N_{0}$ vertices and $\mathcal{E}$ is the set of hyperedges. A hyperedge $e\in\mathcal{E}$ is a subset of $\mathcal{V}$, representing a $\left\lvert e\right\rvert-$body interaction; for example, $e=\{4,7,8\}$ is a 3-body interactions between vertices $4$, 77, and $8$. Since $j$-body interactions correspond to (j1)(j-1)-simplices, following the notation used in the framework of simplicial complexes [^3] When considering simplicial complexes, the order of the interaction is given by the dimension of the space Bianconi (2021). For example, nodes are more formally called 0-simplices, as they have dimension 0, links are $1$-simplices, triangles are $2$-simplices, and so on; hence, a $3$-body interaction is encoded by a $2$-simplex. Bianconi (2021), we refer to $j$-body interactions as (j1)(j-1)-hyperedges and use the subscript or superscript j1j-1 to denote associated structures. In this study, we focus on hypergraphs having only $1$-hyperedges (i.e., $2$-body interactions) and $2$-hyperedges (i.e., $3$-body interactions), and denote the number of $1$- and $2$-hyperedges as $N_{1}$ and $N_{2}$, respectively.
A hypergraph can be represented algebraically as follows. The totality of $2$-hyperedges are represented by the first-order adjacency tensor $A^{(1)}\in\{0,1\}^{N_{0}\times N_{0}}$ (also known as the adjacency matrix), with $A_{jk}=1$ if a $1$-hyperedge exists between vertices $j$ and $k$ (i.e., $\{j,k\}\in\mathcal{E}$), or $A_{jk}=0$ otherwise. Similarly, the totality of $2$-hyperedges are represented by the second-order adjacency tensor $A^{(2)}\in\{0,1\}^{N_{0}\times N_{0}\times N_{0}}$, with $A_{jkl}^{(2)}=1$ if a $2$-hyperedge exists between vertices $j$, $k$, and $l$ (i.e., $\{j,k,l\}\in\mathcal{E}$), or $A_{jkl}^{(2)}=0$ otherwise. $1$-hyperedges can also be represented via the incidence matrix, [^4] Also known as boundary operator in the theory of simplicial complexes Bianconi (2021). denoted by $B^{(1)}\in\{-1,0,1\}^{N_{0}\times N_{1}}$. The $k$-th column of $B^{(1)}$ corresponds to the $k$-th 11-hyperedge, say $\{l,m\}$; this column has $1$ at the $l$-th position, 1-1 at the $m$-th position, and zeros elsewhere. [^5] Formally, assigning $1$ and 1-1 to the vertices of a $1$-hyperedge requires choosing an arbitrary orientation, that is, an ordering of the node pair. This choice does not affect our results.
We say that a hypergraph is $\mathcal{H}$-connected if there exists a path between any pair of vertices using an arbitrary combination of hyperedges of any order. In contrast, we say that a hypergraph is $j$-connected (with $j\in\mathbb{N}$) if there exists a path between any pair of vertices using only $j$-hyperedges; we use in particular the notion of $1$-connectedness (which corresponds to connection via links only). In the literature, a hypergraph that is $\mathcal{H}$-connected is referred to simply as connected Battiston et al. (2020); Nortier et al. (2025); however, we adopt this terminology to clearly distinguish it from the notion of $j$-connectedness, which is useful in the analyses presented in this work. Finally, the $j$-degree (with $j\in\mathbb{N}$) of a vertex $v$, denoted by $d^{(j)}(v)$, is the number of $j$-hyperedges it is a part of. $\langle d^{(j)}\rangle\coloneqq\frac{1}{N_{0}}\sum_{v\in\mathcal{V}}d^{(j)}(v)$ is the average $j$-degree. Note that $\langle d^{(1)}\rangle$ (related to links) corresponds to the usual average degree of a network.
Higher-order Kuramoto models
In the Kuramoto model, higher-order (non-pairwise) interactions naturally emerge from phase reduction Nakao (2016); Pietras and Daffertshofer (2019); Monga et al. (2019). In general, phase reduction theory establishes that it is possible to meaningfully associate a phase oscillator dynamics to a dynamical system in a periodic regime (i.e., whose asymptotic solution is a stable limit cycle) Kuramoto (1984); Winfree (1980); Kuramoto and Nakao (2019). For example, the Kuramoto model itself Kuramoto (1975) can be derived from the complex-valued Stuart-Landau equation Nakao (2014), which can be rewritten as a two-dimensional system admitting a stable limit cycle. In high-dimensional networked systems with weak coupling, a similar phase reduction yields a system of phase oscillators interacting via the underlying network topology Nakao (2016).
The classical Kuramoto model is obtained through a first order approximation with respect to the coupling parameter Kuramoto (1975, 1984). However, when phase reduction is performed using approximations beyond the first order Ashwin and Rodrigues (2016); León and Pazó (2019); Gengel et al. (2020), non-pairwise interactions emerge. As an example, consider three oscillators, $j$, $k$ and $l$, coupled so that both $j$ and $l$ are connected to $k$ but not to each other. While first-order interactions exist only between $j$ and $k$, and $l$ and $k$, a small interaction term appears between $j$ and $l$, when second-order phase reduction is performed [^6] Note that the term higher-order is used ambiguously in the literature, as it can refer both to non-pairwise interactions (many-body coupling) and to higher-order expansions in the coupling parameter. Here, we use higher-order to mean non-pairwise interactions exclusively. Bick et al. (2024).
These non-pairwise terms are typically small corrections, of the order of the squared coupling strength, and thus do not drastically alter the dynamics, but improve the accuracy of the reduced description Sakaguchi (1990); León and Pazó (2019); Mau et al. (2024);
nonetheless, these small higher-order terms can have significant effects near bifurcation points Bick et al. (2024) or in phenomena such as remote synchronization Okuda and Kuramoto (1991); Bergner et al. (2012).
The Kuramoto model with higher-order interactions (i.e., non-pairwise interactions of the same order as pairwise ones) has been extensively studied. However, these higher-order Kuramoto models were typically not derived via phase reduction from high-dimensional oscillators coupled through higher-order interactions. Instead, they were constructed by adding higher-order coupling terms to the classical pairwise Kuramoto model Tanaka and Aoyagi (2011); Skardal and Arenas (2019, 2020); Lucas et al. (2020).
Various forms of higher-order coupling have been proposed in these studies, raising the question of which forms arise naturally from phase reduction. León et al. León et al. (2024) rigorously derived a higher-order Kuramoto model starting from complex Stuart-Landau oscillators coupled via pairwise and higher-order interactions that respect certain symmetry properties. They found that, for the dynamics of the $j$-th oscillator, the naturally emerging $3$-body interaction takes the form [^7] More precisely, the derivation also produces a phase lag, as in the Kuramoto-Sakaguchi model Sakaguchi and Kuramoto (1986). However, since much of the literature focuses on behavior without phase lag, we focus on that setting here. $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ which we refer to as the (1,1,2)(1,1,-2) interaction. This result motivated our choice of the higher-order Kuramoto model, namely,
$\dot{\theta}_{j}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}).$
Note that, unlike the model derived by León et al. León et al. (2024), here the coupling is not global (i.e., all-to-all); hence, the interaction terms are normalized by the average $1$- and $2$-degrees rather than by $N_{0}$ and $N_{0}^{2}$, respectively.
A further generalization of phase reduction for arbitrary hypergraph topologies and coupling functions revealed that also the term $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$—i.e., the (2,1,1)(2,-1,-1) interaction considered [^8] The (2,1,1)(2,-1,-1) interaction studied, e.g., by Skardal and Arenas Skardal and Arenas (2020), allows to apply the Ott-Antonsen reduction Ott and Antonsen (2008), which would not be possible with the (1,1,2)(1,1,-2) interaction. in Skardal and Arenas (2020)—can emerge from phase reduction when the non-reduced system includes certain couplings León et al. (2025). For this reason, we performed our numerical study also on the following higher-order Kuramoto model:
$\dot{\theta}_{j}=\omega_{j}+\frac{K_{1}}{\langle d^{(1)}\rangle}\sum_{k=1}^{N_{0}}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j})\\ +\frac{K_{2}}{2\langle d^{(2)}\rangle}\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\theta_{k}-\theta_{l}-\theta_{j}).$ (2)
The results are in all cases qualitatively analogous to those presented in the Results section, and are reported in the Supplementary Materials.
Synchronization can be measured via the Kuramoto order parameter $R\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{i\theta_{j}}\right\rvert$, where $R\approx 1$ indicates synchronization, and lower values of $R$ correspond to increasing incoherence. To detect two-cluster synchronization, it is possible to use the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$; in that case, $R$ would be low, while $R_{2}$ approaches $1$. In our simulations, we also computed $R_{2}$ and found it generally lower than $R$, indicating that, on average, no significant clustering phenomena occur; these results are shown in the Supplementary Materials.
Lastly, we note that the choice of frequency distribution is arbitrary. We sampled frequencies from uniform distributions, although using unimodal distributions, such as Gaussian or Lorentzian ones, is also valid and commonly used, as it allows some analytical treatment Acebrón et al. (2005). Notably, it is known that a Kuramoto model with unimodal distributions displays the usual second-order transition to synchronization, while a model with uniform frequency distribution Pazó (2005) (or unimodal distributions with a plateau in the center Basnarkov and Urumov (2007)) exhibits a first-order phase transition. However, these differences arise in the thermodynamic limit with global (all-to-all) coupling, which differs significantly from the finite, sparse systems studied here. Indeed, in the Supplementary Materials we report simulations obtained using a Gaussian distribution centered at $0.15$ with standard deviation $0.1$, observing qualitatively similar results to those obtained using the uniform distribution in $[0,0.3]$.
Computationally efficient simulation of the higher-order Kuramoto model
To enable more efficient and scalable numerical simulation of (1), we rewrite it in vector form. Let $z^{(1)},z^{(2)}\in\mathbb{R}^{N_{0}}$ be the vectors collecting all coupling terms associated with $1$- and $2$-hyperedges, respectively, with their $j$-th components given by
$z_{j}^{(1)}\coloneqq\sum_{k=1}^{n}A_{jk}^{(1)}\sin(\theta_{k}-\theta_{j}),\quad z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(\theta_{k}+\theta_{l}-2\theta_{j}).$
For $1$-hyperedges, we first note that $z^{(1)}=B^{(1)}\sin(-(B^{(1)})^{\mathsf{T}}\theta)$, where $B^{(1)}$ is the incidence matrix. Then, for $2$-hyperedges, we define the two auxiliary matrices
$Y^{(2)}_{\mathrm{a}},Y^{(2)}_{\mathrm{s}}\in\mathbb{Z}^{N_{0}\times 3N_{2}}$; the former is used to assemble linear combinations of the phases and the latter to select specific combinations contributing to the dynamics of each oscillator. In $Y^{(2)}_{\mathrm{a}}$, given some $k\in\{1,\dots,N_{2}\}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; the first of these columns has $2$ in position $l$, 1-1 in positions $m$ and $p$, and zeros elsewhere; the second has $2$ in $m$, and 1-1 in $l$ and $p$; the third has $2$ in $p$, and 1-1 in $l$ and $m$. In $Y^{(2)}_{\mathrm{s}}$, the $3$ columns from the (3k2)(3k-2)-th to the $3k$-th also correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; these three columns are different one-hot vectors with $2$ in positions $l$, $m$ and $p$, respectively, and zeros elsewhere. An example set of the auxiliary matrices $Y^{(2)}_{\mathrm{a}}$, $Y^{(2)}_{\mathrm{s}}$ for a hypergraph with $N_{0}=4$ vertices and hyperedges $\{1,2,3\}$, $\{1,3,4\}$ is
$Y^{(2)}_{\mathrm{a}}=\begin{bmatrix}2&-1&-1&&2&-1&-1\\ -1&2&-1&&0&0&0\\ -1&-1&2&&-1&2&-1\\ 0&0&0&&-1&-1&2\end{bmatrix},\quad Y^{(2)}_{\mathrm{s}}=\begin{bmatrix}2&0&0&&2&0&0\\ 0&2&0&&0&0&0\\ 0&0&2&&0&2&0\\ 0&0&0&&0&0&2\end{bmatrix}$.
Then, it is immediate to verify that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\theta)$.
In conclusion, it is possible to rewrite (1) in a vector form as
$\dot{\theta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\theta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\theta\right),$ (3)
where $\theta$ and $\omega$ are $N_{0}$-dimensional column vectors, collecting all $\theta_{j}$’s and $\omega_{j}$’s, respectively. Similar steps can be applied to rewrite (2) in vector form; details are provided in the Supplementary Materials.
Description of the numerical methods
Here, we illustrate the setup of the numerical simulations performed in this work. We integrated the higher-order Kuramoto model using the forward Euler method with time step $0.05$ for $600$ time units. [^9] Initially, results were obtained with a Runge-Kutta IV explicit method and a smaller time step of $0.005$, but since outcomes matched those from the Euler methods we opted for the latter to reduce the high computational cost. Integrating for longer times, e.g., $2000$ time units, also did not affect the results. The order parameter RR—a measure of synchronization—shown in the Results section figures is the time average over the second half of the simulation; time averaging is necessary as, even when starting close to the synchronous state, heterogeneity in frequencies and interactions often prevent the solution from becoming stationary.
Our results represent averages over random hypergraph topologies: for each condition, we sampled $300$ hypergraphs for Figure 3 and $500$ for Figure 5. All simulations were implemented in Matlab The MathWorks Inc. (2024).
Coupling strength analysis
In this analysis, we fix $N_{0}=10$ nodes, $N_{1}=20$ links, and $N_{2}=10$ triangles for the hypergraphs to be generated. Then, we define the coupling strength ranges for pairwise and higher-order interactions, $K_{1}$ and $K_{2}$, both spanning from 0 to $0.6$. A grid of $51\times 51$ points is created over these ranges, and for each grid point, we integrate (1) (or (2), with results shown in the Supplementary Materials [^10] For this case, the range of the coupling strength $K_{2}$ spans from 0 to $1$, as the values of $K_{2}$ maximizing the averaged order parameter $R$ are slightly higher than for the case shown in the Main Text.) over $N_{\mathcal{H}}=300$ randomly generated hypergraph with the prescribed $(N_{0},N_{1},N_{2})$, ensuring either either $\mathcal{H}$-connectivity or $1$-connectivity. For each hypergraph, frequencies $\omega_{j}$ and initial phases $\theta_{0,j}$, for $j=1,\dots,N_{0}$, are drawn uniformly from $[0,0.3]$ and $0,\theta_{0}^{\mathrm{max}}]$, respectively, with $\theta_{0}^{\mathrm{max}}\in\{0.1,2\pi\}$. As noted earlier, the choice of frequency distribution does not qualitatively affect the results. The large number of random hypergraphs, frequency sets, and initial conditions mitigates finite-size effects. Increasing $N_{\mathcal{H}}$ further was tested and showed no significant impact. For each integration, we compute the Kuramoto order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$, averaging their values over the the hypergraph samples at each grid point as described above. When starting from incoherent initial conditions (i.e., $\theta_{0}^{\mathrm{max}}=2\pi$), we also mark, for each $K_{1}$, the value of $K_{2}$ yielding the maximum order parameters.
In Figure 3, we show the order parameter $R$ obtained for the higher-order Kuramoto in (1) starting from both coherent ($\theta_{0}^{\mathrm{max}}=0.1$) and incoherent ($\theta_{0}^{\mathrm{max}}=2\pi$) initial conditions. We also performed the analysis with different numbers of links and triangles, namely $N_{1}=40$, $N_{2}=6$ and $N_{1}=15$, $N_{2}=6$, observing no meaningful difference (results not shown).
In the Supplementary Materials, we show analogous results for the Kuramoto-Daido order parameter $R_{2}$ of the model in (1), for normally distributed frequencies, for larger structures, as well as $R$ and $R_{2}$ for the model in (2). All configurations exhibit the same qualitative behavior shown in the Main Text.
Hyperedge allocation analysis
To determine the optimal allocation of $1$- and $2$-hyperedges for maximizing synchronization, we evaluated the performance of different Kuramoto models with $N_{0}=10$ oscillators, varying hyperedges allocation and coupling strengths. We assigned a fixed cost of $c_{1}=1$ (arbitrary units) to $1$-hyperedges, while $2$-hyperedges had a variable cost $c_{2}\in\{1,3,5\}$. Interpreting a $2$-hyperedge as equivalent to three projected $1$-hyperedges, we consider $c_{2}=3$ to represent cost parity, $c_{2}=1$ to reflect a lower relative cost of $2$-hyperedges, and $c_{2}=5$ to reflect a higher relative cost. The total resource budget was fixed at $J=40$, imposing the constraint
$c_{1}N_{1}+c_{2}N_{2}\leq J.$ (4)
The value of $J$ was chosen to ensure that (i) hypergraphs can always be made $\mathcal{H}$-connected (otherwise synchronization would be impossible) and (ii) the entire resource budget is effectively used, avoiding cases where all possible hyperedges are deployed but unused budget remains. To fulfill these conditions, we note the following facts: with $N_{0}=10$ vertices, the minimum number of $1$-hyperedges for $1$-connectivity is $N_{1}^{\mathrm{min}}=N_{0}-1=9$; the minimum number of $2$-hyperedges for $2$-connectivity is $N_{2}^{\mathrm{min}}=\lceil(N_{0}-1)/2\rceil=5$; the maximum number of distinct $1$-hyperedges is $N_{1}^{\mathrm{max}}=N_{0}(N_{0}-1)/2=45$; the maximum number of distinct $2$-hyperedges is $N_{2}^{\mathrm{max}}=N_{0}(N_{0}-1)(N_{0}-2)/6=120$. Hence, condition (i) requires selecting $J\geq\max\{N_{1}^{\mathrm{min}},5N_{2}^{\mathrm{min}}\}$ (55 is the largest value of $c_{2}$); condition (ii) requires that $J\leq\min\{N_{1}^{\mathrm{max}},1N_{2}^{\mathrm{max}}\}$ (11 is the smallest value of $c_{2}$). This yields an admissible range $J\in[25,45]$, from which we choose $J=40$. To enforce constraint (4), we define $b_{1}\in[0,1]$ as the fraction of resources to be allocated to $1$-hyperedges, and assign $N_{1}=\lfloor Jb_{1}/c_{1}\rfloor$ and $N_{2}=\lfloor J(1-b_{1})/c_{2}\rfloor$.
The experimental conditions were generated by varying coupling strengths $K_{1},K_{2}$ in (1) and the link allocation fraction $b_{1}$. Our goal was to determine the optimal hyperedge allocation for a given physical regime, as characterized by the coupling strengths. To this end, we varied their value while keeping their sum fixed: we selected $K_{1}\in[0,K^{\mathrm{sum}}]$, with $K^{\mathrm{sum}}=0.3$, using $51$ linearly spaced points, and consequently set $K_{2}=K^{\mathrm{sum}}-K_{1}$. The allocation parameter $b_{1}$ was varied over $[0,1]$, also using $51$ linearly spaced points, resulting in $51\times 51=2601$ experimental conditions. For each value of $b_{1}$, we generated $N_{\mathcal{H}}=500$ random hypergraphs and sampled random initial conditions $\theta_{0}$ uniformly from $0,\theta_{0}^{\mathrm{max}}]^{N_{0}}$, as well as natural frequencies uniformly from $0,0.3]^{N_{0}}$, with $\theta_{0}^{\mathrm{max}}$ selected as either $\pi/2$ or $2\pi$. When generating the hypergraphs, the $N_{1}$ and $N_{2}$ hyperedges are selected randomly from all possible ones, each with equal probability. If the resulting hypergraph is not $\mathcal{H}$-connected, it is discarded and a new one is generated in its place, repeating this process until an $\mathcal{H}$-connected hypergraph is obtained.
At this point, a grid of $2601$ experimental conditions are obtained, each with $500$ realizations of hypergraph structure, initial conditions, and natural frequencies. For each realization, we simulate the Kuramoto model (3) and record the average order parameter $R$ in the second half of each simulation. The value of $R$ obtained over all $500$ realizations is averaged to obtain a single value for each experimental condition, thus creating one of the panels in Figure 5. The different panels were created by changing the value of the cost associated with $2$-hyperedges, $c_{2}$, and the range of the initial phase conditions, $\theta_{0}^{\mathrm{max}}$, to produce more or less homogeneous initial states. In total $2601\times 500\times 6=7{,}803{,}000$ simulations were run to realize Figure 5. Also for this analysis, in the Supplementary Materials, we show analogous results for the Kuramoto-Daido order parameter $R_{2}$ of the model in (1), for normally distributed frequencies, for larger structures, as well as $R$ and $R_{2}$ for the model in (2). All configurations exhibit the same qualitative behavior shown here.
Data availability
No datasets were generated or analyzed during the current study. All findings are based on numerical simulations described in the Methods section, which can be reproduced using the details provided therein.
Code availability
The code for the numerical simulations presented in this article is available from the corresponding author upon reasonable request.
Acknowledgments
R.M. and M.C. contributed equally to this work. R.M. acknowledges JSPS KAKENHI 24KF0211 for financial support. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. The authors are grateful to Daniele Proverbio for discussions and feedback on the work. R.M. and M.C. would like to thank Claudio Altafini and all the organizers of the ELLIIT Focus Period on Network Dynamics and Control held at Linköping University in September 2023, where this collaboration started.
Author Contributions
R.M and M.C. conceptualized the work, developed the methodology, carried out the analysis, wrote the manuscript, curated the code, and the visualization. H.N. provided methodological insights. H.N. and M.C. acquired the funding. All authors reviewed and edited the article.
Competing Interests
The authors declare no competing interests.
Supplementary Materials
In the Results section of the Main Text, we presented numerical analyses of the higher-order Kuramoto model with higher-order interaction of the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$ (for the dynamics of the $j$-th oscillator), which we refer to as the (1,1,2)(1,1,-2) interaction; see Eq. (1) in the Main Text. In this Supplementary file, we additionally show results for the model featuring higher-order interactions of the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$—which we refer to as the (2,1,1)(2,-1,-1) interaction (see Eq. (2) in the Main Text)—with a vector form model derived in § Results with a different higher-order interaction. For both types of higher-order interaction, we also provide the value of the Kuramoto-Daido order parameter $R_{2}$ obtained in the simulations (§ Results for the Kuramoto-Daido second order parameter), as well as the simulation results obtained with frequencies distributed normally (§ Results for Gaussian frequency distribution). In § Results for larger hypergraphs, we provide results obtained for larger structures. All results confirm the findings of the Main Text.
All figures presenting simulation results, including those in the Main Text, are summarized in Tables 1 and 2, for the studies on the coupling strength and the hyperedge allocation, respectively.
Figure Discussed in # Nodes Higher-order interaction Metric frequency distribution
Figs. 3, 4 Main Text 10 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM1 Sec. Results with a different higher-order interaction 10 (2,1,1)(2,-1,-1) $R$ uniform
Fig. SM3 Sec. Results for the Kuramoto-Daido second order parameter 10 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM4 Sec. Results for the Kuramoto-Daido second order parameter 10 (2,1,1)(2,-1,-1) $R_{2}$ uniform
Figs. SM7, SM8 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R$ Gaussian
Fig. SM9 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R$ Gaussian
Fig. SM10 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
Fig. SM11 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R_{2}$ Gaussian
Fig. SM16 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R$ uniform, Gaussian
Fig. SM17 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ uniform, Gaussian
Table 1: Review of figures concerning the coupling strength analysis.
Figure Discussed in # Nodes Higher-order interaction Metric frequency distribution
Fig. 5 Main Text 10 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM2 Sec. Results with a different higher-order interaction 10 (2,1,1)(2,-1,-1) $R$ uniform
Fig. SM5 Sec. Results for the Kuramoto-Daido second order parameter 10 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM6 Sec. Results for the Kuramoto-Daido second order parameter 10 (2,1,1)(2,-1,-1) $R_{2}$ uniform
Fig. SM12 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R$ Gaussian
Fig. SM13 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R$ Gaussian
Fig. SM14 Sec. Results for Gaussian frequency distribution 10 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
Fig. SM15 Sec. Results for Gaussian frequency distribution 10 (2,1,1)(2,-1,-1) $R_{2}$ Gaussian
Fig. SM18 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R$ uniform
Fig. SM19 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ uniform
Fig. SM20 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{1}$ Gaussian
Fig. SM21 Sec. Results for larger hypergraphs 100 (1,1,2)(1,1,-2) $R_{2}$ Gaussian
Table 2: Review of figures concerning the hyperedge allocation analysis.
Results with a different higher-order interaction
In the Main Text, we analyzed the higher-order Kuramoto model with (1,1,2)(1,1,-2) higher-order interactions. However, from phase reduction, both second-order pairwise León and Pazó (2019) and first-order many-body León et al. (2025), also the (2,1,1)(2,-1,-1) higher-order interaction can emerge. For this reason,
we complemented the analysis in the Main Text by performing analogous simulations for the Kuramoto model with this alternative higher-order interaction term. To enable more efficient simulations, we reformulated the model in vector form, as detailed below in § Vector form of the higher-order Kuramoto model with the (2,1,1)(2,-1,-1) interaction. The results, shown in Figures SM1, SM2, are qualitatively consistent to those shown in the Main Text, with the following differences. In the coupling strength analysis (Figure SM1), the higher-order interaction coupling strength ($K_{2}$) that maximizes synchronization is slightly larger. In the hyperedge allocation analysis, triangles are less effective at promoting synchronization, which slightly modifies the shape of the black line (indicating optimal allocations) compared to Figure 4 in the Main Text.
Vector form of the higher-order Kuramoto model with the (2,1,1)(2,-1,-1) interaction
Here, we derive a vector form for the Kuramoto model with (2,1,1)(2,-1,-1) higher-order interactions. As in the Main Text, the model can be written in vector form as
$\dot{\theta}=\omega+\frac{K_{1}}{\langle d^{(1)}\rangle}B_{1}\sin\left(-(B_{1})^{\mathsf{T}}\theta\right)+\frac{K_{2}}{2\langle d^{(2)}\rangle}Y^{(2)}_{\mathrm{s}}\sin\left(-(Y^{(2)}_{\mathrm{a}})^{\mathsf{T}}\theta\right),$ (SM1)
but with different matrices $Y^{(2)}_{\mathrm{a}}$ and $Y^{(2)}_{\mathrm{s}}$. Namely, let $z^{(2)}\in\mathbb{R}^{N_{0}}$ be the vector collecting all coupling terms of the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$ Skardal and Arenas (2020) associated with $2$-hyperedges, with the $j$-th component given by
$z_{j}^{(2)}\coloneqq\sum_{k=1}^{N_{0}}\sum_{l=1}^{N_{0}}A_{jkl}^{(2)}\sin(2\theta_{k}-\theta_{l}-\theta_{j}).$
Define the two auxiliary matrices $Y^{(2)}_{\mathrm{a}},Y^{(2)}_{\mathrm{s}}\in\mathbb{Z}^{N_{0}\times 6N_{2}}$. In $Y^{(2)}_{\mathrm{a}}$, given some $k\in\{1,\dots,N_{2}\}$, the $6$ columns from the (6k5)(6k-5)-th to the $6k$-th correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; the first of these columns has $1$ in positions $l$ and $p$, 2-2 in position $m$, and zeros elsewhere; the second has $1$ in $l,m$, and 2-2 in $p$; the third has $1$ in $m,p$, and 2-2 in $l$; the fourth has $1$ in $m,l$, and 2-2 in $p$; the fifth has $1$ in $p,m$, and 2-2 in $l$; the sixth has $1$ in $p,l$, and 2-2 in $m$. In $Y^{(2)}_{\mathrm{s}}$, the $6$ columns from the (6k5)(6k-5)-th to the $6k$-th also correspond to the $k$-th 22-hyperedge, say $\{l,m,p\}$; these six columns are different one-hot vectors with $1$ in positions $l$, $l$, $m$, $m$, $p$, and $p$, respectively, and zeros elsewhere. An example set of the auxiliary matrices $Y^{(2)}_{\mathrm{a}}$, $Y^{(2)}_{\mathrm{s}}$ for a hypergraph with $N_{0}=4$ vertices and hyperedges $\{1,2,3\}$, $\{1,3,4\}$ is
[…]
Then, it holds that $z^{(2)}=Y^{(2)}_{\mathrm{s}}\sin(-(Y^{(2)}_{\mathrm{b}})^{\mathsf{T}}\theta)$, which confirms the form in Eq. (SM1).
Results for the Kuramoto-Daido second order parameter
In Figures SM3, SM4, SM10, SM11, SM17 (for the coupling strength analysis) and SM5, SM6, SM14, SM15, SM19, SM21 (for the hyperedge allocation analysis), we show the value of the Kuramoto-Daido order parameter $R_{2}\coloneqq\left\lvert\frac{1}{N_{0}}\sum_{j=1}^{N_{0}}e^{2i\theta_{j}}\right\rvert$ obtained for the simulations considered, with both kinds of higher-order interactions. Comparing these results with those in the Main Text and in this Supplementary Materials, we observe that $R_{2}$ is large in the same regions where $R$ is large, indicating clustering phenomena are absent, on average.
Results for Gaussian frequency distribution
In Figures SM7, SM8, SM9, SM10, SM11, SM16, SM17 (for the coupling strength analysis) and SM12, SM13, SM14, SM15, SM20, SM21 (for the hyperedge allocation analysis), we show the order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$ obtained for all the simulations considered, in the case that, for each hypergraph realization, oscillator frequencies $\omega_{j}$ are drawn randomly from a Gaussian distribution with mean $0.15$ and standard deviation $0.1$.
The results are qualitatively analogous to those reported in the Main Text, where a uniform frequency distribution was considered instead.
Results for larger hypergraphs
In Figures SM16, SM17 (for the coupling strength analysis) and SM18, SM19, SM20, SM21 (for the hyperedge allocation analysis), we show the value of the order parameter $R$ and the Kuramoto-Daido order parameter $R_{2}$ obtained for simulations of larger hypergraphs of $N_{0}=100$, with both uniform and normal frequency distributions.
To illustrate the choice of the simulation parameters, we recall the expressions for (i) the minimum number $N_{1}^{\mathrm{min}}$ of links required for $1$-connectivity, (ii) the minimum number $N_{2}^{\mathrm{min}}$ of triangles required for $2$-connectivity , (iii) the maximum number $N_{1}^{\mathrm{max}}$ of distinct links, and (iv) the maximum number $N_{2}^{\mathrm{max}}$ of distinct triangles. For a structure of $N_{0}$ nodes, these are
$N_{1}^{\mathrm{min}}(N_{0})=N_{0}-1,\qquad N_{2}^{\mathrm{min}}(N_{0})=\left\lceil\frac{N_{0}-1}{2}\right\rceil,\qquad N_{1}^{\mathrm{max}}(N_{0})=\frac{N_{0}(N_{0}-1)}{2},\qquad N_{2}^{\mathrm{max}}(N_{0})=\frac{N_{0}(N_{0}-1)(N_{0}-2)}{6}.$
For the coupling strength analysis, we set the numbers of links $N_{1}$ and triangles $N_{2}$ to as to preserve the same interaction densities used for the smaller hypergraphs ($N_{0}=10$). Preserving link and triangle densities requires keeping the ratios $N_{1}/N_{1}^{\mathrm{max}}$ and $N_{2}/N_{2}^{\mathrm{max}}$ constant. For $N_{0}=10$, we have $N_{1}^{\mathrm{max}}(10)=45$ and $N_{2}^{\mathrm{max}}(10)=120$, and we set $N_{1}=20$ and $N_{2}=10$. Then, for $N_{0}=100$, we compute $N_{1}^{\mathrm{max}}(100)=4950$ and $N_{2}^{\mathrm{max}}(100)=161{,}700$; thus, matching the previous densities yields $N_{1}=2200$ and $N_{2}=13{,}475$.
Owing to the very high computational cost associated with such a large number of triangles, in Figs. SM16 and SM17 we adopted a coarser $4\times 7$ parameter grid and a reduced number of hypergraph realizations $N_{\mathcal{H}}=100$. In this case, the higher-order Kuramoto model was integrated for 1000 time units (instead of $600$), as convergence required longer times (convergence was further verified by extending the simulations up to $3000$ time units, with no qualitative changes observed). Despite the coarser grid, the results confirm the trends observed for smaller structures.
For choosing the budget value $J$ for the hyperedge allocation analysis, we proceeded as follows. Let $c_{1}^{\mathrm{min}}$ (resp, $c_{2}^{\mathrm{min}}$) be the minimum cost associated to links (resp. triangles), and let $c_{1}^{\mathrm{max}}$ (resp, $c_{2}^{\mathrm{max}}$) denote the corresponding maximum cost. In our study, $c_{1}^{\mathrm{min}}=c_{1}^{\mathrm{max}}=1$, $c_{2}^{\mathrm{min}}=1$, and $c_{2}^{\mathrm{max}}=5$. The budget $J$ was then chosen to satisfy two conditions: (i) $J\geq\max\{c_{1}^{\mathrm{max}}N_{1}^{\mathrm{min}},c_{2}^{\mathrm{max}}N_{2}^{\mathrm{min}}\}$, ensuring $\mathcal{H}$-connected hypergraphs could always be generated; (ii) $J\leq\min\{c_{1}^{\mathrm{min}}N_{1}^{\mathrm{max}},c_{2}^{\mathrm{min}}N_{2}^{\mathrm{max}}\}$, so that no portion of the budget is wasted (see Methods for details). For $N_{0}=100$, we have $N_{1}^{\mathrm{min}}(100)=99$, $N_{2}^{\mathrm{min}}(100)=50$, $N_{1}^{\mathrm{max}}(100)=4950$, and $N_{2}^{\mathrm{max}}(100)=161{,}700$. Hence, conditions (i) and (ii) yield $J\geq\max\{1\cdot 99,5\cdot 50\}=250$ and $J\leq\max\{1\cdot 4950,1\cdot 161{,}700\}=4950$. This defines the admissible range $J\in[250,4950]$, from which we selected $J=600$. In this case, we used a $31\times 31$ parameter grid and $N_{\mathcal{H}}=100$ hypergraph realizations. The qualitative findings reported in the Main Text are confirmed also in this setting.
Figure SM1: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM2: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM3: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$ (see Methods). Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM4: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ uniformly distributed in $[0,0.3]$, and initial phases $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM5: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM6: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM7: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM8: System response to different pairwise ($K_{1}$) and higher-order ($K_{2}$) coupling strengths, for initially incoherent $1$-connected hypergraphs from Fig. SM7. The top-center panel replicates the bottom right panel of Fig. SM7 for comparison (color is the mean order parameter $R$). The top-left (resp. top-right) panel shows the mean value of $R$ (blue line) and its standard deviation (shaded area) as a function of the higher-order coupling strength $K_{2}$ (resp. $K_{1}$) while keeping the pairwise interaction strength $K_{1}$ (resp. $K_{2}$) fixed. The values of $K_{1}$, $K_{2}$ explored are also indicated in the top-center panel by a horizontal (resp. vertical) black line. Pink (resp. brown) triangles pointing upward/downward (resp. leftward/rightward) mark representative pairs $(K_{1},K_{2})$. The bottom panels display the distributions (integral normalized to 11) of $R$ across the $N_{\mathcal{H}}=300$ realizations of hypergraph, initial conditions, and frequencies.
Figure SM9: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R$.
Figure SM10: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R_{2}$.
Figure SM11: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=300$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. The grid $K_{1},K_{2}$ is $51\times 51$. Each hypergraph has $N_{0}=10$ nodes, $N_{1}=20$ links and $N_{2}=10$ triangles, with oscillator frequencies $\omega_{j}$ distributed normally with average $0.15$ and standard deviation $0.1$, and initial phases distributed uniformly in $0,\theta_{0}^{\mathrm{max}}]$. Left panels show the cases of initial phases close to synchronization, i.e., $\theta_{0}^{\mathrm{max}}=0.1$; right panels show the case of incoherent initial states, i.e., $\theta_{0}^{\mathrm{max}}=2\pi$. On the right panels, the black line indicates, for each $K_{1}$, the $K_{2}$ yielding the maximum $R_{2}$.
Figure SM12: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM13: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM14: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM15: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(2\theta_{k}-\theta_{l}-\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $51\times 51$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=500$ random hypergraphs were generated, with $N_{0}=10$ nodes, frequencies distributed normally with average $0.15$ and standard deviation $0.1$ and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panels, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panels. The relative link cost is fixed at $c_{1}=1$; triangle costs $c_{2}$ are $1$ in top panels, 33 in middle panels, and $5$ in bottom panels. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM16: Average Kuramoto order parameter $R$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=100$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $4\times 7$. Each hypergraph has $N_{0}=100$ nodes, $N_{1}=2200$ links and $N_{2}=13{,}475$ triangles. On the left panel, oscillator frequencies $\omega_{j}$ are uniformly distributed in $[0,0.3]$, while on the right panel they are distributed normally with mean $0.15$ and standard deviation $0.1$. Initial phases are drawn uniformly in $0,\theta_{0}^{\mathrm{max}}=2\pi]$.
Figure SM17: Average Kuramoto-Daido order parameter $R_{2}$ as a function of the coupling strengths of pairwise ($K_{1}$) and higher-order interactions ($K_{2}$), computed over $N_{\mathcal{H}}=100$ randomly generated hypergraphs. The higher-order interactions in the Kuramoto model have the form $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. The grid $K_{1},K_{2}$ is $4\times 7$. Each hypergraph has $N_{0}=100$ nodes, $N_{1}=2200$ links and $N_{2}=13{,}475$ triangles. On the left panel, oscillator frequencies $\omega_{j}$ are uniformly distributed in $[0,0.3]$, while on the right panel they are distributed normally with mean $0.15$ and standard deviation $0.1$. Initial phases are drawn uniformly in $0,\theta_{0}^{\mathrm{max}}=2\pi]$.
Figure SM18: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM19: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies drawn uniformly at random from $[0,0.3]$, and initial phases from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM20: Average Kuramoto order parameter $R$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies distributed normally with mean $0.15$ and standard deviation $0.1$, and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).
Figure SM21: Average Kuramoto-Daido order parameter $R_{2}$ for different combinations of links and triangles, using the higher-order coupling $\sin(\theta_{k}+\theta_{l}-2\theta_{j})$. $K_{1}$ and $K_{2}$ are the coupling strengths of pairwise and $3$-body interactions, respectively. The grid $(b_{1},K_{1})$ is $31\times 31$. For each value of the link allocation fraction $b_{1}$, $N_{\mathcal{H}}=100$ random hypergraphs were generated, with $N_{0}=100$ nodes, frequencies distributed normally with mean $0.15$ and standard deviation $0.1$, and initial phases drawn uniformly from $0,\theta_{0}^{\mathrm{max}}]$$\theta_{0}^{\mathrm{max}}=\frac{\pi}{2}$ in the left panel, and $\theta_{0}^{\mathrm{max}}=2\pi$ in the right panel. The relative link cost is fixed at $c_{1}=1$, while triangle cost is $c_{2}=3$. For each $K_{1}$ (y-axis), the black line marks the $b_{1}$ (x-axis) yielding the highest $R$ (color).